OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line that passes through the given point . y=3=8/3 (x+2) ; (-2,3)

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

I think you made a mistake in your equation. There are two equal signs.

OpenStudy (anonymous):

But what you need to do is find the slope of that equation, take its negative reciprocal(this will be the slope of the perpendicular line) and substitute this slope into the equation of a line with that point. ie. Find m when the equation is put into the form: y = mx + b Take -1 divided by m for the new slope. To find b of the new line, use b = y - mx